OpenStudy (anonymous):

A car with mass 1137 kg, moving at 17 m / s, strikes a(n) 2540 kg car at rest. If the two cars stick together, with what speed do they move?

OpenStudy (loser66):

elastic collision? hey, this is physics question, why do you post here?

OpenStudy (anonymous):

didnt know this wasnt for physics . oops

OpenStudy (loser66):

but I know @IDKwut knows how to solve, let him help

OpenStudy (loser66):

where is he???

OpenStudy (idkwut):

m1 v1 = m1 V1 + m2 V2 = (m1 + m2) V Plug and solve for V.

OpenStudy (loser66):

yup and v2 =0 so that it 's easy to solve, right?

OpenStudy (anonymous):

in not to sure what to plug in where, this is new to me

OpenStudy (idkwut):

O. I'll give you another hint: V = m1/(m1 + m2) v1 Where m1 = mass 1, m2 = mass 2 and v1 = final velocity. Try it out. :D

OpenStudy (anonymous):

is the answer 0.007087

OpenStudy (idkwut):

Please tell me how you got that.

OpenStudy (anonymous):

m1-1137 m2-17 v1-2540

OpenStudy (idkwut):

17 is not the second mass...

OpenStudy (anonymous):

2540?

OpenStudy (idkwut):

Well, that's a given since it is in your question post. m1 = 1137, m2 = 2540, and velocity is measured in m/s which is 17 m/s, now plug into the formula I gave you. :)

OpenStudy (anonymous):

thank you so much.

OpenStudy (anonymous):

43197 is the answer?

OpenStudy (idkwut):

Tell me how you got that now. I want to see how you work the problem out.

OpenStudy (anonymous):

\[V = \frac{ m1 }{ (m1+m2) } \times v1\] Order of operations? Shouldn't it be \[V = \frac{ m1 \times v1 }{ (m1 + m2)}\]

OpenStudy (anonymous):

Follow order of operations do not multiply the bottom and divide

OpenStudy (idkwut):

Don't give her/him the answer. It's pretty simple work to do.

OpenStudy (anonymous):

They are the same thing as long as you follow the operations

OpenStudy (anonymous):

Was not going to

OpenStudy (anonymous):

Re-stated the formula that was given above and gave a guideline of how order of operations matters here especially if they are plugging numbers into a calculator

OpenStudy (anonymous):

5.257

OpenStudy (idkwut):

Good job! :)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

are you able to help with this one A(n) 1090 N man stands in the middle of a frozen pond of radius 6 . 09 m. He is un- able to get to the other side because of a lack of friction between his shoes and the ice. To overcome this difficulty, he throws his 1 . 05 kg physics textbook horizontally toward the north shore, at a speed of 4 . 53 m / s. How long does it take him to reach the south shore?